### v1

# deep neural network STFT short-time Fourier transform

Abstract
We present ReverbFX, a new room impulse response (RIR) dataset designed for singing voice dereverberation research. Unlike existing datasets based on real recorded RIRs, ReverbFX features a diverse collection of RIRs captured from various reverb audio effect plugins commonly used in music production. We conduct comprehensive experiments using the proposed dataset to benchmark the challenge of dereverberation of singing voice recordings affected by artificial reverbs. We train two state-of-the-art generative models using ReverbFX and demonstrate that models trained with plugin-derived RIRs outperform those trained on realistic RIRs in artificial reverb scenarios.
11footnotetext: Authors contributed equally to this work.
1 Introduction
[…]
RIR play a crucial role in audio signal processing, as they characterize how sound propagates and decays in a given environment [1]. To obtain room impulse responses that characterize natural reverberation, they can either be measured in actual rooms or simulated using established acoustic models [2]. These natural RIRs reflect the physical properties and geometries of actual spaces, yielding reverberation with consistent time decay, energy distribution, and frequency response characteristics. Accordingly, most RIR datasets and dereverberation benchmarks are based on this natural room acoustics paradigm, aiming to improve technologies such as speech recognition, hearing aids, and audio enhancement in real-world environments [3].
However, there is a discrepancy between natural reverberation and reverb effects used in modern music production. In music production, often artificial RIRs are employed using digital reverb audio effect plugins. Unlike natural reverberation, artificial reverbs are unconstrained by the laws of physics and can exhibit a wide range of exotic behaviors: extremely long or modulated decays, non-monotonic frequency responses, non-linear characteristics, and even deliberately unnatural acoustics. These reverberation effects are part of the creative palette, shaping the aesthetics of contemporary music. As a consequence, dereverberation approaches trained and evaluated exclusively on natural RIR datasets may not generalize well to these artificial reverberation scenarios.
To address this gap, we introduce ReverbFX, a novel dataset of RIRs entirely generated from a diverse set of professional reverb audio effect plugins. Unlike existing RIR datasets, ReverbFX captures the diversity and complexity of artificial reverbs frequently used in music production. This unique dataset enables, for the first time, the study and benchmarking of dereverberation models specifically for these artificial environments.
We use ReverbFX to create a new benchmark, SingingReverbFX, for evaluating singing voice dereverberation under artificial reverberation conditions. Using this data, we train two state-of-the-art generative models [4, 5] and demonstrate that, when trained on SingingReverbFX, these models can effectively dereverberate singing voices recorded with artificial reverbs, significantly outperforming counterparts trained solely on natural RIRs (i.e., in an off-domain setting). Our experiments establish ReverbFX as an essential resource for advancing dereverberation research beyond natural acoustics and towards applications in music production, where artificial reverbs predominate.
In summary, our contributions are three-fold:
  1. 1.
    We motivate and construct ReverbFX, the first dataset focused on artificial reverberation from audio effect plugins.
  2. 2.
    We systematically compare two generative dereverberation models trained on artificial versus natural reverberation.
  3. 3.
    We establish a new benchmark for singing voice dereverberation in artificial reverb scenarios, paving the way for more robust and musically relevant audio enhancement technologies.
2 Background
Reverberation, or reverb for short, is a phenomenon that occurs in enclosed spaces when sound signals reflect off walls and other surfaces [3]. In music production, reverb is one of the most commonly used effects, applied to almost every track in a mix. This is especially true for vocals, where reverb is used to create tension, add texture, or simply achieve a more natural and immersive sound. Reverb can make vocals sound wider and significantly influence the emotional response of listeners; for example, music with long, dense reverb often evokes an epic or dreamlike atmosphere.
While reverb is a powerful tool for enhancing emotion and depth in music, it can severely degrade the quality and intelligibility of speech and vocals. This is particularly problematic for individuals using hearing aids, for whom reverb can create substantial comprehension difficulties [6, 7]. In music production, it is usually preferable to apply artificial reverb to non-reverberant (dry) vocals, since this allows precise parameter control, which requires the removal of existing natural or artificial reverb. Non-reverberant vocals are also preferred for further music production tasks such as remixing a song, as they provide greater flexibility for creative processing.
Removing reverb from a signal, especially from vocals or speech, is a challenging problem in the field of signal processing. Singing voice dereverberation is conceptually similar to speech dereverberation, and similar methods can be used to address both tasks. However, despite their similarities, speech and singing vocals have distinct characteristics, particularly in terms of pitch, power, and phoneme duration.
Early approaches to vocal and music dereverberation primarily relied on signal processing techniques such as linear prediction [8, 9, 10, 11], or source models designed to capture the harmonic structure of musical signals [12, 13]. These methods typically assumed a structured representation of the source and exploited music’s inherent periodicity and tonal characteristics to mitigate reverberation effects.
Recent research in machine learning has introduced a variety of new approaches, including end-to-end deep neural networks [14]. Although initially proposed for reverb conversion, such models can also be applied to dereverberation. However, these supervised methods tend to learn the characteristics of specific reverbs present in the training data, leading to reduced performance when encountering unseen reverb types, a limitation given the wide range of possible reverberation scenarios in real-world applications. To address this challenge, more recent work has increasingly focused on generative approaches, such as diffusion models, which can be supervised or unsupervised and aim to achieve robust generalization across diverse reverb conditions.
Feedback Delay Networks
One of the foundational concepts behind algorithmic reverb plugins was introduced by Stautner and Puckette in 1982 [15]. Their approach feeds multiple copies of the input signal into a network of delay lines with varying lengths. The outputs from these delays are then weighted and recursively routed back into the system through a feedback delay matrix, laying the groundwork for what would later be formalized as feedback delay networks [16]. This architecture allows the signal to circulate within the network, mimicking the complex reflections in natural acoustic environments. The feedback matrix is typically unitary to ensure filter stability, resulting in a lossless FDN. The reverb tail’s envelope is generally shaped using a combination of gain-based decay and low-pass filtering on the delay lines, simulating the gradual energy loss, especially of high frequencies, that occurs with multiple reflections in real spaces.
Beyond the feedback process, some algorithms also include a feedforward path that routes the input signal through the delay lines once before reaching the output. This component helps approximate early reflections and provides additional dry/wet balance control.
Figure 1: Histogram of RT60 values for ReverbFX dataset, showing the distribution of reverberation times among the collected RIRs.
3 ReverbFX Dataset
We release the ReverbFX dataset [^1] https://sp-uhh.github.io/reverbfx, comprising 1,846 RIRs, with RT60 times ranging from 0.31 to 52.08 seconds. A histogram of RT60 times is shown in Figure 1. All RIRs were generated using four publicly available audio plugins, with further details provided in Section 3.1. To measure each RIR, a Dirac impulse was used as the input signal; this was created from a WAV file in which the first sample is set to 1 and all subsequent samples are set to 0. The reverb effects were applied using DawDreamer [17], a Python API that provides essential DAW functionalities, including integration of Steinberg VST plugins.
For each plugin, we utilized both factory and custom presets. To further increase variation in the resulting audio effects, we generated 14 randomized versions of each preset by modifying two randomly selected parameters from a predefined subset (e.g., delay time and low cut). This procedure results in a total of 15 impulse responses per preset. For all configurations, the dry/wet parameter was explicitly set to 1 to ensure a fully wet signal.
A rigorous validation procedure was employed to filter out anomalous or corrupted data to guarantee the quality and reliability of the RIRs. Each RIR underwent the following checks:
  • Numerical Integrity: Verification for invalid numerical values such as NaNs or infinite entries.
  • DC Offset: Ensuring the mean value is close to zero (threshold: $10^{-4}$)
  • Energy Threshold: Discarding RIRs with an energy below $10^{-16}$
  • Length Validation: Excluding empty or zero-length RIRs.
  • RT60 Consistency: Retaining only those RIRs with a computed RT60 in the plausible range of 0.1 to 100 seconds.
  • Hilbert Envelope: Evaluating the amplitude envelope via the Hilbert transform to exclude RIRs with excessively short decay times.
  • Early Energy Concentration: Assessing the proportion of energy in the initial segment to ensure it is not overly concentrated at the start.
Dataset Duration # Singers # Files Song type Language $f_{s}$ Bit depth
(Hours) (kHz) (bit)
OpenSinger [18] $51.9$ $25$ m, $41$ f $39116$ Traditional and Pop Songs Chinese $24$ $16$
M4Singer [19] $29.8$ $10$ m, $10$ f $20696$ Pop Songs Mandarin $44.1$ $24$
CSD [20] $4.7$ $1$ f $1810$ Children Songs Korean, English $44.1$ $16$
PJS [21] $0.5$ $1$ m $195$ Children and Anime Songs Japanese $48$ $24$
Opencpop [22] $5.2$ $1$ f $2517$ Chinese Pop Songs Chinese $44.1$ $24$
NUS-48E [23] $1.9$ $6$ m, $6$ f $703$ Traditional and Pop Songs English $44.1$ $16$
NHSS [24] (Test) $4.8$ $5$ m, $5$ f $490$ Pop Songs English $48$ $16$
Table 1: Overview of vocal datasets used for training and the test. The number of speakers is indicated by their gender, male (m) or female (f), and $f_{s}$ denotes the sample rate. The number of files reflects only those included in the dataset and does not count files discarded during the generation process.
3.1 Reverb Audio Effect Plugins
Protoverb (v1.0.1)
The Protoverb plugin is proprietary software distributed by u-he Heckmann Audio GmbH [^2] https://u-he.com and licensed for free, non-commercial use. Contrary to traditional algorithmic reverbs that aim to minimize resonances, Protoverb builds up as many room resonances as possible, modeling the body of air in a room [25]. It is based on a delay network, whose structure is controlled by a text-based randomization feature, allowing users to generate unique reverb settings. For our study, we created $200$ presets with different random strings in the reverb section to capture a diverse set of RIRs
SkyNet (v0.4)
SkyNet Reverb by Shift Line  [^3] https://shift-line.com is proprietary software offered under a Pay-What-You-Want model.
The SkyNet plugin is based on the two reverb algorithms SkyNet-1 and SkyNet-2 from the Astronaut space reverb pedal, and is designed to create expansive soundscapes with long decay times and diffuse sound [26]. The plugin offers a parameter to control the ratio between the two algorithms.
TAL-Reverb-4 (v4.0.4)
The Tal-Reverb-4 Plugin is proprietary software distributed by TAL Software GmbH [^4] https://tal-software.com. It is based on a modulated delay network and is designed to evoke the character of 1980s digital reverbs [27]. The algorithm produces a high-quality, plate-style reverb with a diffuse and modulated character, making it particularly suitable for vocals and synthetic sounds.
Valhalla Supermassive (v3.0.0)
The Valhalla Supermassive plugin is proprietary software distributed by Valhalla DSP [^5] https://valhalladsp.com and offered as freeware. Designed for extreme and lush reverb and delay effects, it implements a modulated FDN architecture optimized for massive spatialization, long decay times, and evolving textures [28]. The plugin offers a range of algorithm, each with unique temporal and spectral characteristics, making it particularly suited for ambient sound design and experimental processing.
3.2 License
We obtained explicit permission from the plugin developers to generate and publish RIRs derived from the plugins for non-commercial research purposes. The released dataset does not contain any part of the original software. ReverbFX is released under the Creative Commons Attribution-NonCommercial 4.0 International License (CC BY-NC 4.0). It may be used, shared, and adapted for non-commercial research and educational purposes only. Commercial use is strictly prohibited. Proper attribution must be given to the dataset authors and the original plugin developers.
4 SingingReverbFX Benchmark
We introduce SingingReverbFX, a singing voice dereverberation benchmark focused on restoring reverberant vocals processed with audio reverb effects commonly used in music production. To this end, we convolve singing voice data with our new ReverbFX dataset of RIRs captured from reverb effect plugins.
4.1 Data
Inspired by Lemercier et al. [29], we utilize a variety of singing voice datasets, including OpenSinger [18], M4Singer [19], CSD [20], PJS [21], OpenCpop [22], NUS-48E [23], and NHSS [24] These datasets feature studio-quality recordings with balanced gender and phoneme distributions and include vocals in multiple languages. The combined vocal data amounts to a total of 94 hours. An overview of the datasets and their key characteristics is provided in Table 1.
To prepare data for training, validation, and the test, we partition the ReverbFX dataset using stratified sampling based on RT60 bins, ensuring that each subset is representative of the overall distribution of reverberation times. The data is split into training, validation, and test sets with a ratio of 1446:200:200. The vocal training set includes all recordings except those set aside for validation and testing. For validation, we select 980 gender-balanced files from the OpenSinger dataset. The entire NHSS dataset is used exclusively for testing.
For every split, each vocal recording is paired with a randomly selected RIR from the corresponding subset. To ensure compatibility for convolution, vocal files are first resampled to 48 kHz. Each RIR is truncated at its maximum index to remove trailing silence or noise, then normalized, converted to mono, and convolved with the vocal signal.
After convolution, we apply dry/wet mixing, a standard parameter in digital reverb plugins, which defines the balance between the original (dry) and reverberated (wet) vocals. At each iteration, the mixing ratio $\alpha$ is randomly sampled from a uniform distribution $\alpha\in[0.1,1]$, resulting in
$y(t)=(1-\alpha)x(t)+\alpha(x(t)\ast h(t)),$ (1)
where $y(t)$ is the final mixture, $x(t)$ is the dry vocal, and $h(t)$ is the RIR.
Our approach, convolving vocals with RIRs, leverages the principles of linearity and time invariance (LTI), meaning that only linear, time-invariant reverb characteristics can be captured. Some reverb plugins, however, include time-varying or nonlinear elements (e.g., Chorus or Pitch Modulation) that cannot be replicated through RIR convolution. Nevertheless, using RIRs offers a straightforward way to construct the dataset while still modeling essential reverb parameters, such as delay time, intensity, and absorption characteristics.
POLQA PESQ ESTOI MAE $\downarrow$ OVRL REVERB
Reverberant $1.89\pm 1.01$ $1.61\pm 0.76$ $0.41\pm 0.29$ $0.14\pm 0.12$ $1.81\pm 0.69$ $3.11\pm 0.58$
SGMSE+ [4] (mism.) $2.53\pm 1.31$ $2.23\pm 1.06$ $0.56\pm 0.27$ $0.14\pm 0.10$ $2.72\pm 0.47$ $4.22\pm 0.39$
SGMSE+ [4] $\mathbf{3.12\pm 1.33}$ $\mathbf{2.63\pm 1.10}$ $\mathbf{0.64\pm 0.24}$ $\mathbf{0.08\pm 0.06}$ $\mathbf{2.99\pm 0.32}$ $\mathbf{4.30\pm 0.38}$
Schrödinger Bridge [5] $2.90\pm 1.25$ $2.37\pm 1.05$ $0.61\pm 0.24$ $0.13\pm 0.09$ $2.85\pm 0.42$ $4.12\pm 0.46$
Table 2: Singing voice dereverberation results on SingingReverbFX using standard speech enhancement metrics. Values indicate mean and standard deviation. Higher values indicate better performance for all metrics except MAE.
4.2 Baselines
All models are trained under identical conditions using two NVIDIA RTX A6000 graphics processing units, with a batch size of 6. Training is stopped after 350 k steps. The following paragraphs provide details about the specific models used in our experiments.
Score-based Generative Models
Richter et al. [4] introduced SGMSE+, a score-based generative model for speech enhancement and dereverberation. In a supervised setting, SGMSE+ learns clean speech posteriors conditioned on noisy or reverberant inputs. The model operates in the complex short-time Fourier transform (STFT) domain, utilizing a diffusion process and a neural network to estimate the score function, which guides the reverse generative process.
In a follow-up study [30], the authors adapted the model for 48 kHz audio and conducted an extensive hyperparameter search to optimize the noise schedule and amplitude compression of the input representation. We adopt these optimal hyperparameters to train SGMSE+ on SingingReverbFX using the official implementation. [^6] https://github.com/sp-uhh/sgmse As an ablation, we also train SGMSE+ on natural RIRs, using the datasets employed in [30], to illustrate the domain shift between natural and artificial reverberation.
Schrödinger Bridge
Jucic et al. [5] extended SGMSE+ by employing a tractable Schrödinger bridge as the diffusion process for speech enhancement and dereverberation. The Schrödinger bridge transforms the input speech distribution into the clean speech distribution by formulating and solving an optimal transport problem. Richter et al. [31, 32] later adapted this model for 48 kHz audio and integrated it into their code repository, which we use for training the Schrödinger bridge on SingingReverbFX. We utilize the stochastic differential equation (SDE) sampler, as it has been shown to provide superior performance for dereverberation tasks [5].
4.3 Metrics
Currently, there are no metrics specifically designed for singing voice dereverberation. Therefore, we adopt a set of established metrics from speech enhancement to evaluate both perceptual quality, intelligibility, and signal fidelity:
  • POLQA [33]: An intrusive, model-based standard (ITU-T P.863) for perceptual audio quality, considered a successor to PESQ and widely used in speech evaluation.
  • PESQ [34]: An intrusive metric that estimates perceived speech quality on a scale from 1 to 4.5, originally developed for telephony.
  • ESTOI [35]: Measures intelligibility by comparing short-time spectral features between reference and output, giving a score from 0 (unintelligible) to 1 (perfect).
  • MAE: Mean absolute error in the STFT domain, providing a direct, objective comparison between estimated and reference signals.
  • OVRL [36]: A non-intrusive neural network-based estimator that predicts mean opinion score (MOS) for overall quality.
  • REVERB [36]: A non-intrusive neural network-based estimator that predicts MOS for perceived reverberation.
Together, these metrics provide a comprehensive evaluation framework to assess both the perceptual and objective quality of dereverberated singing voice signals.
5 Results
In Table 2, we present the singing voice dereverberation results obtained using SingingReverbFX. It can be seen that all methods yield improvements over the reverberant test set across all metrics. Furthermore, training SGMSE+ on natural reverb (with mismatched RIRs) leads to a performance decrease compared to training SGMSE+ with in-domain data from the SingingReverbFX training set. The Schrödinger bridge also demonstrates strong performance, but is slightly outperformed by SGMSE+. This may be attributed to the extensive hyperparameter tuning applied to the 48 kHz model in [30].
6 Conclusions
In this work, we introduced ReverbFX, the first dataset dedicated to artificial reverberation effects commonly used in music production. Through comprehensive benchmarking with state-of-the-art generative models, we demonstrated that models trained on plugin-derived RIRs significantly outperform those trained on traditional, natural RIRs when faced with dereverberation of singing voices under artificial reverb conditions. Our results highlight the importance of domain-specific data for dereverberation tasks and establish ReverbFX and the Singing-ReverbFX benchmark as vital resources for developing and evaluating audio enhancement models tailored to modern music production environments.